OpenStudy (anonymous):

Antonio is seven less than four times his daughter's age. Antonio is 37; how old is his daughter? 11 12 18 6

OpenStudy (goformit100):

18

OpenStudy (anonymous):

11 would be the right answer

OpenStudy (anonymous):

since 11*4 = 44 - 7 = 37

OpenStudy (anonymous):

4d - 7 = 37 Thats the equation he used